OpenStudy (anonymous):

Prove: Δ/s-a = s*tan a/2

OpenStudy (anonymous):

\[\frac{ Δ }{ s-a } = s \tan \frac{ \alpha }{ 2 }\]

OpenStudy (anonymous):

using R.H.S ??

OpenStudy (anonymous):

\[\tan \frac{ \alpha }{ 2 } = \sqrt{\frac{ (s-b)(s-c) }{ s(s-a) }}\]

OpenStudy (anonymous):

and \[\Delta = \sqrt{s(s-a)(s-b)(s-c)}\]

hartnn (hartnn):

then its direct. \(\dfrac{\Delta}{s-a} =\dfrac{ \sqrt{s(s-a)(s-b)(s-c)}}{s-a}=\dfrac{ \sqrt{s(s-b)(s-c)}}{\sqrt{s-a}}\)

hartnn (hartnn):

\(s\tan \frac{ \alpha }{ 2 } =s \sqrt{\frac{ (s-b)(s-c) }{ s(s-a) }}= \sqrt{\frac{s (s-b)(s-c) }{ (s-a) }}\)

hartnn (hartnn):

LHS = RHS.

OpenStudy (anonymous):

we have to prove it using a single side

hartnn (hartnn):

okk.

OpenStudy (anonymous):

i think we should choose R.H.S

hartnn (hartnn):

any1 can be chosen., \(\dfrac{\Delta}{s-a} =\dfrac{ \sqrt{s(s-a)(s-b)(s-c)}}{s-a}\\ =\dfrac{ \sqrt{s(s-b)(s-c)}}{\sqrt{s-a}} \\=\sqrt{\frac{s (s-b)(s-c) }{ (s-a) }} \\ =s \sqrt{\frac{ (s-b)(s-c) }{ s(s-a) }} \\ = s\tan \frac{ \alpha }{ 2 } \)

OpenStudy (anonymous):

WHAT HAPPENED IN THE SECOND STEP WITH s-a ???

hartnn (hartnn):

\(\dfrac{\sqrt{s-a}}{s-a}=\dfrac{1}{\sqrt{s-a}}\)